INAF LBC+LUCIFER queue observing, 2010 December 03 - 04

Observer: K.Boutsia, M. Fumana,V. Testa
Telescope Operator: S. Allanson
Support Astronomer: M. Pedani
Instrument: LUCIFER

Summary

Some clouds, spectroscopic weather, start with Lucifer and probably go on with it all night long
The clouds persisted throughout the night
Some problems with the position of the N3.75 camera (~30min lost) and the AGW (~15min lost)
The first half of the night had a seeing <1.0'', but it got better during the second half with seeing <0.7''

Data Summary

Project Objects Filter ExpTime Notes
Pentericci cluste1HK 200JH OrderSep 4hr ARC+FLAT , the seeing degraded from 0.5'' to 1.0'' during the execution, possible clouds
Cresci HDF 210zJHK+H 3hr20min ARC+FLAT + Tell, seeing <0.7''

Details

00:15 - Open Dome, cyrrhus everywhere, Luci in the beam
  • Wait for Sunset

00:45 - Problems in getting Lucifer's camera in

01:05 - Acquisition for Cresci DSF223 (campoA) 0001-0003
  • problems with the centering of N3.75

01:10 - Acquisition DFS223 (campoA) again after doing INITIALIZE 004-007
  • MOS not well centered
  • we decide to abort and readjust manually the camera N3.75

01:39 - Camera re-centered

01:40 - Slew to Pentericci cluster1HK for MOS acquisition 0008-0011
  • We can not see the pivot stars in the slit - probably because it is cloudy
  • we decide to proceed with the spectroscopy

01:57 - Science spectra for cluster1HK 0012- 0015
  • OB aborted after the 4th offset because we can see no spectra
  • either MOS is not well centered or it is because of the clouds

02:25 - We try to recenter manually the N3.75 camera
  • sky images to check the centering of the N3.75 0016-0022

03:13 - Acquisition of cluster1HK 0023-0028
  • we had to do the acquisition twice because at some point the N1.8 camera was inserted for no apparent reason

03:32 - Start science MOS spectroscopy for cluster1HK (Pentericci) 1stHR 0029-0040
  • still no sign of a spectrum in the difference images
  • probably due to clouds

04:38 - acquisition for cluster1HK 0041-0044
  • we repeat the acquisition just to be sure for the MOS positioning since there is no sign of spectra

04:51 - Start science MOS spectroscopy for cluster1HK (Pentericci) 2ndHR 0045-0056

05:57 - Science MOS spectroscopy for cluster1HK (Pentericci) 3rdHR 0057-0068
  • we have applied an offset Y=-1.5 to recenter the mask without acquisition

07:03 - Commit offset, Science MOS spectroscopy for cluster1HK (Pentericci) 4thHR 0069-0080
  • Sky conditions variable, clouds keep passing over, sometimes cannot see objects
  • Sky lines shift from one spectrum to the other
  • Guide star FWHM degraded from 0.5 to 1.0 since the beginning of Pentericci's observations

08:08 - Calibration Unit IN
  • ARC 0081-0084
  • FLAT 0085-0088

08:16 - Acquisition Telluric for Cresci HDFN 0089

08:20 - Spectra for Telluric HDFN in H band 0090-0097

08:30 - MOS acquisition for HDFN
  • after 10min it does not manage to collimate
  • probably guiding probe is vibrating because of the low elevation
  • pupils also look strange
  • clear all optics - collimate at a near by Perssons and send the preset again

08:52 - MOS acquisition for HDFN again 0098-0101
  • after the re-initialization of AGW guide star looks fine

09:00 - Science spectra MOS for HDFN
  • lost guide star - preset lost - need to restart PCS

09:03 -we have to repeat MOS acquisition HDFN 0102-0105

09:16 - Science spectra MOS for HDFN in H 0106-0117
  • guide star looks fine - seeing 0.7''-0.8''

10:23 - Science spectra MOS for HDFN in H 0118-0129
  • Give Y+1.5 offset to go back to original pointing

11:30 - Science spectra MOS for HDFN in H 0130-0141
  • Give Y+1.5 offset to go back to original pointing

12:37 - Science spectra MOS for HDFN in H 0142-0145
  • Give Y+1.5 offset to go back to original pointing
  • Abort script after 4th iteration because sky is too bright
  • Seeing ~0.6''

13:00 - Calib Unit IN
  • ARC H 0146-0149
  • FLAT H 0150-0155

13:10 - Tell acquisition for HDFN in H 0156-0157

13:14 - Tell MOS spectrum for HDFN in H 0158-0165
  • after the end of the spectroscopy, we took some images of the Tell through the slit, without the grating to verify the position 0166-0169

13:22 - END OF NIGHT

-- VincenzoTesta - 04 Dec 2010
Topic revision: r5 - 15 Mar 2012, VincenzoTesta
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback